OpenStudy (natalie1234):

HELP! WILL GIVE MEDAL AND BECOME A FAN!!!! y varies inversely with x and y = 8 when x = 2. Which is the inverse variation equation for the relationship?

OpenStudy (joemarket):

The inverse variation equation is when x=8, y=2.

OpenStudy (natalie1234):

my options are y = 16x y = 16/x y = 2x y = x + 2